12) A student takes a 10-question multiple-choice test by guessing. Each question has 4 choices. Use the binomial distribution to compute the following probabilities. a. b. c. The student gets at most 2 correct. The student gets exactly 3 correct. The student gets at least 7 correct.
10) Compute the mean, variance, and standard deviation for the probability distribution listed below. 4 P(X) 1/3 1/8 1/8 1/4 1/6
Answer #1

12)here for binomial distribution parameter n=10 and p=1/4 (probability of getting a question correct)

a)P(X<=2) =P(X=0)+P(X=1)+P(X=2)=10C0(1/4)0(3/4)10+10C1(1/4)1(3/4)9+10C2(1/4)2(3/4)8=0.5256

b)P(X=3)=+10C3(1/4)3(3/4)7 =0.2503

c)

P(X>=7)=P(X=7)+P(X=8)+P(X=9)+P(X=10)

=10C7(1/4)7(3/4)3+10C8(1/4)8(3/4)2+10C9(1/4)9(3/4)1 +10C10(1/4)10(3/4)0 =0.0035

10)

x f(x) xP(x) x2P(x)
1 1/3 0.333 0.333
2 1/8 0.250 0.500
3 1/8 0.375 1.125
4 1/4 1.000 4.000
5 1/6 0.833 4.167
total 2.792 10.125
E(x) =μ= ΣxP(x) = 2.7917
E(x2) = Σx2P(x) = 10.1250
Var(x)=σ2 = E(x2)-(E(x))2= 2.3316
std deviation=         σ= √σ2 = 1.5270

from above mean =2.7917

variance= 2.3316

std deviation=1.5270